OpenStudy (anonymous):

Which number is greater than 7.3 × 10^8?

OpenStudy (anonymous):

@iGreen

geerky42 (geerky42):

Well, there is infinitely many numbers that is greater than 7.3 × 10^8

geerky42 (geerky42):

You could easily say like 7.4 × 10^8

geerky42 (geerky42):

or 8 × 10^8

OpenStudy (igreen):

Yeah, that's true. First evaluate 10^8.

OpenStudy (igreen):

10^ 8 = = 100000000 8 x 100000000 = 800000000 So any number greater than 800000000 will be your answer.

OpenStudy (anonymous):

1.135 x 10^8 3.3 x 10^9 8.46 x 10^7 9.2 x 10^6

geerky42 (geerky42):

Which do you think? @Future_UsArmy_MP

OpenStudy (igreen):

Future, this is obvious. Look at the exponents. 10^6 = Six zeroes after '1'. 10^7 = Seven zeroes after '1'. So which do you think is the answer now? @Future_UsArmy_MP